OpenStudy (anonymous):

brad wants to buy flowers for his friend with $45. the daisies are $1 each and the roses are $2 each. he buys 3 more daisies than roses. How much did the roses cost?

OpenStudy (mathstudent55):

He buys d daisies and r roses. His expense is 1d + 2r The total expense is 45, so our first equation is: d + 2r = 45 Then we are also told about the number of each flower bought: "he buys 3 more daisies than roses" That translates into: d = r + 3 You need to solve the system of equations below: d + 2r = 45 d = r + 3

OpenStudy (anonymous):

how would i solve the system of equations

OpenStudy (zzr0ck3r):

d + 2r = 45 d = r + 3 so plug r+3 in for d in the first equation r+3+2r=45 3r=42 r=?

OpenStudy (anonymous):

r=14

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

so the roses cost 14

OpenStudy (zzr0ck3r):

yes and d = 3+r = 3+ 14 = 17

OpenStudy (anonymous):

but the roses dont cost 14 the answer was wrong

OpenStudy (mathstudent55):

r = number of roses. The solution to the equation shows that 14 roses were bought. Each cost $2. The money spent on roses is 14 x $2 = $28